OpenStudy (anonymous):

3-4 (z+1 )= 3z -(2+z)

OpenStudy (anonymous):

3-4z-4=3z-2-z -1-4z=2z-2 -4z=2z-1 -6z=-1 6z=1 z=1/6

OpenStudy (anonymous):

Just so you know, you can plug exactly what you posted into wolfram alpha and it will give you the solution much more quickly. http://www.wolframalpha.com/